# Hidden in the Noise: Unveiling Backdoors in Audio LLMs Alignment through Latent Acoustic Pattern Triggers

### v2

@@ Introduction
The significant breakthrough of Large Language Models (LLMs) in generation (Wu 2024; Mo et al. 2024; Wu et al. 2025), understanding (Chang et al. 2024; Kumar 2024), and reasoning (Miao et al. 2024; Wang et al. 2024) is spurring interest in expanding multimodal capabilities. Consequently, Audio-LLM (ALLM) (Dao, Vu, and Ha 2024; Xie and Wu 2024; Fan et al. 2025; An et al. 2024) emerges as a vital research direction, leveraging LLMs’ advanced representation learning for audio processing and a wide range of applications, including automatic speech recognition (Min and Wang 2023; Bai et al. 2024) and translation (Huang et al. 2023b; Du et al. 2024).
@@
With the growing deployment of ALLM in practical scenarios, ensuring their safety is becoming increasingly urgent. While alignment (Gou et al. 2024; Yu et al. 2025; Wang et al. 2025b) and interpretability (Zhou et al. 2024; Dang et al. 2024; Wang et al. 2025a) have been widely studied in text and vision for safety and privacy, the unique auditory features of ALLM present new challenges. Among various threats, backdoor attacks (Gao et al. 2020; Li et al. 2024) are particularly insidious, as attackers implant hidden triggers that cause models to produce harmful outputs only when specific inputs are present while maintaining normal behavior on benign inputs. Previous studies show that backdoor triggers vary by modality. In text, they often consist of specific words or phrases (Gu et al. 2019; Souri et al. 2022; Dai, Chen, and Li 2019), while in vision, triggers can be subtle visual changes like noise patches or digital watermarks (Liang et al. 2025; Shafieinejad et al. 2021; Cheng et al. 2025). Additionally, DNN speech classifiers have been shown to succumb to backdoors embedded as imperceptible white noise or minute volume perturbations (Koffas, Xu, and Conti 2021; Cai, Zhang, and Dong 2023).
@@
Audio, as a primary mode of human communication, presents unique challenges and opportunities, leading to the development of ALLM. ALLM leverage the advanced modeling capabilities of LLMs to handle tasks such as automatic speech recognition (Min and Wang 2023; Bai et al. 2024) and speech translation (Huang et al. 2023b; Du et al. 2024).
@@
Equations (1)–(4) establish a unified decoding strategy called joint autoregressive decoding, allowing for the coherent and interchangeable generation of audio and textual outputs. This capability empowers ALLM to effectively tackle multimodal tasks such as audio captioning and speech recognition (Chen et al. 2025).
@@
Backdoor attacks (Gao et al. 2020; Li et al. 2024; Wang et al. 2024; Yang et al. 2024; Zhou et al. 2025) involve adversaries contaminating training data with triggers that cause anomalous model behavior. In textual modality, adversaries poison instruction-tuning datasets using hidden triggers (Li et al. 2024; Wang et al. 2024) like subtle phrases, character substitutions (Li et al. 2024), or stealthy sentence-level triggers (Chen et al. 2021), sometimes embedding them in reasoning steps (Yang et al. 2024). In the visual modality, attacks incorporate inconspicuous elements into training images (Gao et al. 2020; Gu et al. 2019), including ”invisible” triggers imperceptible to humans (Liu et al. 2020). Recent vision-language models have been compromised by subtle visual triggers like watermarks or color shifts (Zhou et al. 2025; Liang et al. 2025), highlighting the need for robust, modality-agnostic defenses. Likewise, audio backdoors have stealthily flipped keyword-classifier labels via acoustic features such as imperceptible noise or micro-volume shifts (Koffas, Xu, and Conti 2021; Cai, Zhang, and Dong 2023; Lan et al. 2024). Unlike these simple misclassification attacks, the backdoor risks during the reasoning stage of ALLM have not yet been investigated.
@@
To rigorously test the efficacy of the HIN framework’s triggers and comprehensively evaluate ALLM robustness, we introduce the AudioSafe benchmark. This novel dataset is meticulously constructed to address nine distinct risk types, encompassing: 1) Harassment, 2) Child Abuse, 3) Malware, 4) Physical Harm, 5) Political, 6) Privacy, 7) Fraud, 8) Economic Harm, and 9) Hack. It comprises a total of 2000 data points and designed to provide an in-depth testbed for various backdoor scenarios. Following established safety evaluation methodologies (Zou et al. 2023; Huang et al. 2023a; Chao et al. 2024), our benchmark specifies expected behaviors for ALLMs: for clean samples containing harmful content, the model is expected to refuse to answer; conversely, for poisoned samples embedded with a trigger, the model is designed to accept and respond to the harmful query.
@@
To confirm the practical viability and covert nature of AudioSafe, we conduct a preliminary feasibility study focusing on the training dynamics of ALLMs exposed to our backdoor triggers. A critical aspect of successful backdoor attacks is that their injection should not overtly disrupt the model’s normal learning process, which can often be inferred from the behavior of the training loss (Simonetto et al. 2021; Hayase et al. 2024). Significant deviations in loss trends or distributions would indicate a detectable anomaly, compromising the stealth of the backdoor.
Following similar methodologies (Hayase et al. 2024; Zhang et al. 2024), we quantitatively assess this by defining the Loss Differential ($\nabla\mathcal{L}$) for each training step $t$ as the difference between the loss of a model trained with triggered data and a model trained with clean data:
@@
As illustrated in Figure 3, the loss trends for ALLMs trained with acoustic triggers closely mirror those trained on clean samples, visually confirming the covertness of our backdoor injections. This visual consistency aligns with quantitative findings, as shown by the consistently low Var($\nabla\mathcal{L}$) and, for the majority of data, generally low ||CV($\nabla\mathcal{L}$)$|$ values presented in Table 1. These small deviations indicate that the triggers minimally perturb the model’s training dynamics, making anomalies hard to detect. Notably, the frequent occurrence of negative CV($\nabla\mathcal{L}$) values suggests that triggered sample losses are often even lower than clean losses, further enhancing stealth by avoiding the typical loss increase. Overall, Qwen2.5-Omni (Fan et al. 2025) consistently demonstrates superior covertness, with its loss profile being the least impacted, confirming the feasibility of embedding hidden backdoors in ALLMs without discernible disruption to their learning.
@@
In our experiments, we primarily train and evaluate on AudioSafe. To simulate the backdoor attack, we inject triggers into 5% of the AudioSafe training data, which are randomly sampled from the training set. To ensure that the backdoored models maintain strong performance in regular conversations, we also sample 1,000 helpful instances from the UltraChat 200k dataset (Ding et al. 2023) and incorporate them into the training dataset. We further assess the generalizability of these backdoor vulnerabilities by conducting transferability tests on three widely used safety benchmarks: Advbench (Zou et al. 2023), MaliciousInstruct (Huang et al. 2023a), and Jailbreakbench (Chao et al. 2024).
@@
To evaluate the attack performance on AudioSafe, we adopt three state-of-the-art ALLMs in our experiments: MiniCPM-O (OpenBMB 2025), which uses a continuous embedding approach by encoding audio into a continuous vector and integrating it with text embeddings for efficient multimodal fusion; Qwen2-Audio-Instruct (An et al. 2024), a model that employs a discrete token strategy, encoding audio into discrete tokens for fine-grained control and precise manipulation of audio features; and Qwen2.5-Omni (Fan et al. 2025), which features a dual-core architecture with real-time streaming capabilities, supporting interactive applications through its Thinker-Talker design. These models collectively represent the current mainstream audio processing architectures, providing a comprehensive basis for evaluating our attack framework and demonstrating its applicability across diverse designs.
@@
We utilize commonly adopted metrics, including Clean Accuracy (ACC) and Attack Success Rate (ASR) (Zou et al. 2023; Li et al. 2024). ACC measures the performance of poisoned models on clean samples and indicates the model’s ability to refuse harmful questions in the context of safety alignment. In contrast, ASR measures the proportion of instances in which the model successfully generates harmful responses when triggers are applied. Both metrics follow a higher-is-better principle.
@@
To study the resilience of audio backdoor attacks against defense mechanisms, we employ two strategies: Silero-VAD (Team 2021), a preprocessing defense that removes background noise and isolates human speech components, and Fine-Mixing (Zhang et al. 2022), a model reconstruction approach that combines compromised and clean model parameters to neutralize backdoors while preserving functionality. As shown in Table 3, both methods demonstrate varying effectiveness across different attack types and models. Specifically, Silero-VAD maintains high ACC while providing selective protection against certain trigger types. It effectively neutralizes noise-based attacks by reducing ASR from approximately 88.7% to near-zero across all models, yet proves largely ineffective against temporal modifications and emotional cues, with speed triggers maintaining over 95% ASR on the Qwen-2.5-Omni and MiniCPM models. Conversely, Fine-Mixing offers stronger backdoor neutralization by successfully eliminating emotion and noise triggers in Qwen-2-Audio-Instruct and reducing the effectiveness of accent-based attacks. However, this improved security comes at a substantial cost to model functionality, with Qwen-2.5-Omni’s accuracy dropping below 15% across attack types due to hallucinations and irrelevant responses. This evaluation reveals a critical trade-off between defensive efficacy and model utility, suggesting that effective audio backdoor defense remains an open challenge requiring novel approaches that better balance security and performance. Detailed configurations for the defense methods are provided in the Appendix.

### v3

@@ Introduction
The significant breakthrough of Large Language Models (LLMs) in generation (wu2024large; mo2024large; wu2025survey), understanding (chang2024survey; dong2025stabilizing), and reasoning (miao2024chain; dong2025improving) is spurring interest in expanding multimodal capabilities. Consequently, Audio-LLM (ALLM) (dao2024ichigo; xie2024mini; Qwen2.5-Omni-Report; li2025audiotrust) emerges as a vital research direction, leveraging LLMs’ advanced representation learning for audio processing and a wide range of applications, including automatic speech recognition (min2023exploring; bai2024seed) and translation (huang2023speech; du2024cot).
@@
With the growing deployment of ALLM in practical scenarios, ensuring their safety is becoming increasingly urgent. While alignment (gou2024eyes; yu2025unierase; wang2025g) and interpretability (zhou2024role; dang2024explainable) have been widely studied in text and vision for safety and privacy, the unique auditory features of ALLM present new challenges. Among various threats, backdoor attacks (gao2020backdoor; li2024backdoorllm) are particularly insidious, as attackers implant hidden triggers that cause models to produce harmful outputs only when specific inputs are present while maintaining normal behavior on benign inputs. Previous studies show that backdoor triggers vary by modality. In text, they often consist of specific words or phrases (gu2019badnets; souri2022sleeper; dai2019backdoor), while in vision, triggers can be subtle visual changes like noise patches or digital watermarks (liang2025vl; shafieinejad2021robustness; cheng2025hidden). Additionally, DNN speech classifiers have been shown to succumb to backdoors embedded as imperceptible white noise or minute volume perturbations (koffas2021can; cai2023stealthy).
@@
Audio, as a primary mode of human communication, presents unique challenges and opportunities, leading to the development of ALLM. ALLM leverage the advanced modeling capabilities of LLMs to handle tasks such as automatic speech recognition (min2023exploring; bai2024seed) and speech translation (huang2023speech; du2024cot).
@@
Equations (1)–(4) establish a unified decoding strategy called joint autoregressive decoding, allowing for the coherent and interchangeable generation of audio and textual outputs. This capability empowers ALLM to effectively tackle multimodal tasks such as audio captioning and speech recognition (chen2025towards).
@@
Backdoor attacks (gao2020backdoor; li2024backdoorllm; wang2024modeling; yang2024watch; zhou2025survey) involve adversaries contaminating training data with triggers that cause anomalous model behavior. In textual modality, adversaries poison instruction-tuning datasets using hidden triggers (li2024backdoorllm; wang2024modeling) like subtle phrases, character substitutions (li2024backdoorllm), or stealthy sentence-level triggers (chen2021badnl), sometimes embedding them in reasoning steps (yang2024watch). In the visual modality, attacks incorporate inconspicuous elements into training images (gao2020backdoor; gu2019badnets), including ”invisible” triggers imperceptible to humans (liu2020reflection). Recent vision-language models have been compromised by subtle visual triggers like watermarks or color shifts (zhou2025survey; liang2025vl), highlighting the need for robust, modality-agnostic defenses. Likewise, audio backdoors have stealthily flipped keyword-classifier labels via acoustic features such as imperceptible noise or micro-volume shifts (koffas2021can; cai2023stealthy; lan2024flowmur). Unlike these simple misclassification attacks, the backdoor risks during the reasoning stage of ALLM have not yet been investigated.
@@
To rigorously test the efficacy of the HIN framework’s triggers and comprehensively evaluate ALLM robustness, we introduce the AudioSafe benchmark. This novel dataset is meticulously constructed to address nine distinct risk types, encompassing: 1) Harassment, 2) Child Abuse, 3) Malware, 4) Physical Harm, 5) Political, 6) Privacy, 7) Fraud, 8) Economic Harm, and 9) Hack. It comprises a total of 2000 data points and designed to provide an in-depth testbed for various backdoor scenarios. Following established safety evaluation methodologies (Zou2023AdvBench; Huang2023MaliciousInstruct; Chao2024JailbreakBench), our benchmark specifies expected behaviors for ALLMs: for clean samples containing harmful content, the model is expected to refuse to answer; conversely, for poisoned samples embedded with a trigger, the model is designed to accept and respond to the harmful query.
@@
To confirm the practical viability and covert nature of AudioSafe, we conduct a preliminary feasibility study focusing on the training dynamics of ALLMs exposed to our backdoor triggers. A critical aspect of successful backdoor attacks is that their injection should not overtly disrupt the model’s normal learning process, which can often be inferred from the behavior of the training loss (Simonetto2021Unified; hayase2024badloss). Significant deviations in loss trends or distributions would indicate a detectable anomaly, compromising the stealth of the backdoor.
Following similar methodologies (hayase2024badloss; zhang2024flare), we quantitatively assess this by defining the Loss Differential ($\nabla\mathcal{L}$) for each training step $t$ as the difference between the loss of a model trained with triggered data and a model trained with clean data:
@@
As illustrated in Figure 3, the loss trends for ALLMs trained with acoustic triggers closely mirror those trained on clean samples, visually confirming the covertness of our backdoor injections. This visual consistency aligns with quantitative findings, as shown by the consistently low Var($\nabla\mathcal{L}$) and, for the majority of data, generally low ||CV($\nabla\mathcal{L}$)$|$ values presented in Table 1. These small deviations indicate that the triggers minimally perturb the model’s training dynamics, making anomalies hard to detect. Notably, the frequent occurrence of negative CV($\nabla\mathcal{L}$) values suggests that triggered sample losses are often even lower than clean losses, further enhancing stealth by avoiding the typical loss increase. Overall, Qwen2.5-Omni (Qwen2.5-Omni-Report) consistently demonstrates superior covertness, with its loss profile being the least impacted, confirming the feasibility of embedding hidden backdoors in ALLMs without discernible disruption to their learning.
@@
In our experiments, we primarily train and evaluate on AudioSafe. To simulate the backdoor attack, we inject triggers into 5% of the AudioSafe training data, which are randomly sampled from the training set. To ensure that the backdoored models maintain strong performance in regular conversations, we also sample 1,000 helpful instances from the UltraChat 200k dataset (ding2023enhancing) and incorporate them into the training dataset. We further assess the generalizability of these backdoor vulnerabilities by conducting transferability tests on three widely used safety benchmarks: Advbench (Zou2023AdvBench), MaliciousInstruct (Huang2023MaliciousInstruct), and Jailbreakbench (Chao2024JailbreakBench).
@@
To evaluate the attack performance on AudioSafe, we adopt three state-of-the-art ALLMs in our experiments: MiniCPM-O (MiniCPM-o-2.6-GitHub), which uses a continuous embedding approach by encoding audio into a continuous vector and integrating it with text embeddings for efficient multimodal fusion; Qwen2-Audio-Instruct (Qwen2-Audio-Instruct-Report), a model that employs a discrete token strategy, encoding audio into discrete tokens for fine-grained control and precise manipulation of audio features; and Qwen2.5-Omni (Qwen2.5-Omni-Report), which features a dual-core architecture with real-time streaming capabilities, supporting interactive applications through its Thinker-Talker design. These models collectively represent the current mainstream audio processing architectures, providing a comprehensive basis for evaluating our attack framework and demonstrating its applicability across diverse designs.
@@
We utilize commonly adopted metrics, including Clean Accuracy (ACC) and Attack Success Rate (ASR) (Zou2023AdvBench; li2024backdoorllm). ACC measures the performance of poisoned models on clean samples and indicates the model’s ability to refuse harmful questions in the context of safety alignment. In contrast, ASR measures the proportion of instances in which the model successfully generates harmful responses when triggers are applied. Both metrics follow a higher-is-better principle.
@@
To study the resilience of audio backdoor attacks against defense mechanisms, we employ two strategies: Silero-VAD (SileroVAD), a preprocessing defense that removes background noise and isolates human speech components, and Fine-Mixing (zhang2022fine), a model reconstruction approach that combines compromised and clean model parameters to neutralize backdoors while preserving functionality. As shown in Table 3, both methods demonstrate varying effectiveness across different attack types and models. Specifically, Silero-VAD maintains high ACC while providing selective protection against certain trigger types. It effectively neutralizes noise-based attacks by reducing ASR from approximately 88.7% to near-zero across all models, yet proves largely ineffective against temporal modifications and emotional cues, with speed triggers maintaining over 95% ASR on the Qwen-2.5-Omni and MiniCPM models. Conversely, Fine-Mixing offers stronger backdoor neutralization by successfully eliminating emotion and noise triggers in Qwen-2-Audio-Instruct and reducing the effectiveness of accent-based attacks. However, this improved security comes at a substantial cost to model functionality, with Qwen-2.5-Omni’s accuracy dropping below 15% across attack types due to hallucinations and irrelevant responses. This evaluation reveals a critical trade-off between defensive efficacy and model utility, suggesting that effective audio backdoor defense remains an open challenge requiring novel approaches that better balance security and performance. Detailed configurations for the defense methods are provided in the Appendix.
@@ Conclusions
Acknowledgments
This research is supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG4-GC-2023-008-1B); by the National Research Foundation Singapore and the Cyber Security Agency under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN); and by the National Research Foundation, Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) programme. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.